### v1

# No.2 Linggong Road, Dalian, Liaoning, 116024, P.R.China

Abstract
Neural networks have emerged as a powerful paradigm for tasks in high energy physics, yet their opaque training process renders them as a black box. In contrast, the traditional cut flow method offers simplicity and interpretability but demands human effort to identify optimal boundaries. To merge the strengths of both approaches, we propose the Learnable Cut Flow (LCF), a neural network that transforms the traditional cut selection into a fully differentiable, data-driven process. LCF implements two cut strategies—parallel, where observable distributions are treated independently, and sequential, where prior cuts shape subsequent ones—to flexibly determine optimal boundaries. Building on this, we introduce the Learnable Importance, a metric that quantifies feature importance and adjusts their contributions to the loss accordingly, offering model-driven insights unlike ad-hoc metrics.  To ensure differentiability, a modified loss function replaces hard cuts with mask operations, preserving data shape throughout the training process. LCF is tested on six varied mock datasets and a realistic diboson vs. QCD dataset. Results demonstrate that LCF (1) accurately learns cut boundaries across typical feature distributions in both parallel and sequential strategies, (2) assigns higher importance to discriminative features with minimal overlap, (3) handles redundant or correlated features robustly, and (4) performs effectively in real-world scenarios. In diboson dataset, LCF initially underperforms boosted decision trees and multiplayer perceptrons when using all observables. However, pruning less critical features—guided by learned importance—boosts its performance to match or exceed these baselines. LCF bridges the gap between traditional cut flow method and modern black-box neural networks, delivering actionable insights into the training process and feature importance.
1 Introduction
At the Large Hadron Collider (LHC), vast data from high-energy collisions are collected, allowing physicists to search for signals of Beyond the Standard Model (BSM) physics against overwhelming background noise. As a common tool, the cut flow (cut-based) method applies a series of manually defined cuts in sequence to filter rare signal events. While straightforward in eliminating background and preserving signals, it struggles with rigidity as the complexity and number of observables grow.
In contrast to the cut flow method, machine learning is increasingly adopted across the field, proving effective in diverse applications: tagging jets with deep networks and jet charge 1511.05190 ; 1908.08256 ; 2012.08526 ; 2305.04372 , speeding up shower simulations with generative models 2210.06204 ; 2301.08128 ; 2305.17169 ; 2312.00042 ; 2401.02248 ; 2407.16704 ; 2410.21611 , enhancing unfolding for full-event reconstruction 2104.03036 ; 2105.09923 ; 2112.08180 ; 2203.16722 ; 2404.14332 , anomaly detection in complex LHC data 2305.15179 ; 2309.13111 ; 2310.06897 ; 2310.13057 ; 2311.17162 ; 2312.10009 , and suppressing pileup with attention-based methods 2107.02779 ; 2203.15823 .
While machine learning excels in high-energy physics, its opacity has spurred efforts to align it with physical intuition and traditional methods. One approach embeds symmetry principles into neural networks, yielding architectures like Lorentz-equivariant autoencoders 2212.07347 , PELICAN networks 2211.00454 , energy flow networks 1810.05165 ; 2012.00964 , particle cloud jet tagging 1902.08570 , and others 2203.06153 ; 2212.00832 ; 2411.00446 . Another bridges neural networks with high-level observables, with ref. 1704.08249 leveraging ML to quantify jet information and saturate feature space, ref. 1710.01305 deriving novel jet observables, and ref. 1902.07180 automating observable construction. Additionally, researchers map low-level features to high-level observables 2010.11998 and craft interpretable anomaly detectors for jet substructure 2203.01343 , extracting discriminative insights in a human-readable form. For a comprehensive overview, see ref. 2102.02770 .
To overcome the opacity of neural networks while retaining the clarity of traditional cut-based analysis, we introduce the Learnable Cut Flow (LCF). It is a neural network that fully emulates human cut-searching operations, ensuring its training process is entirely transparent and understandable. At bottom, we transform cuts into differentiable operations, such that training a neuron is equivalent to searching for an optimal cut. This match reflects LCF’s transparency and interpretability. For scenarios where signals lie in the middle or at both edges, we employ mask operations to split observable distributions into two parts around a specified center. With such learnable cuts, boundaries are automatically identified after training.
Since observables vary in their separation power, contributing unevenly to classification, we introduce Learnable Importance. It is a set of softmax-normalized trainable weights that favors stronger observables while suppressing weaker ones. Unlike other ad-hoc metrics, it not only reflects the contribution fraction but also provides a robust way for selecting key observables to build the final cut flow.
To build the LCF model, we stack a series of layers. It starts with data preprocessing for standard normalization. Next comes the learnable importance layer. A split layer then separates data, passing each observable to its corresponding learnable cut. These cuts, matching the number of observables, process the data. Finally, a concatenation layer collects all cut results. During training, this structure learns the full cut flow. At inference, we select results based on importance—retaining cuts from high-importance observables and discarding the rest.
We train the LCF model by rethinking human cut-searching strategies: a parallel approach searches each raw observable distribution independently, while a sequential approach searches the distribution altered by prior cuts. In the parallel strategy, we sum the loss across all cuts. For the sequential strategy, we modify the loss function to incorporate prior cuts. Here, mask operations apply binary weights to filter out contributions from already-cut events. This ensures training aligns with LCF’s transparent design.
We demonstrate LCF’s effectiveness on dedicated mock datasets and a realistic diboson vs. QCD dataset. Results show it accurately learns cut boundaries in both strategies, prioritizes discriminative features with minimal overlap, handles redundant or correlated features robustly, and performs well in real-world cases.
The paper is structured as follows: section 2 explores the methodological foundations, section 3 details the experimental setups, section 4 presents results and analysis, and section 5 concludes the study.
2 Methodology
2.1 Learnable cuts
In high-energy physics, traditional cuts flows depend heavily on human effort. First, physicists select relevant observables. Next, each observable is visualized as a one-dimensional distribution. Then, boundaries are chosen by inspecting these distributions to separate signal from background events.
To formalize cuts, consider a dataset with $N$ event samples ($i$ indexes events) and $F$ observables ($j$ indexes observables). A traditional cut on $x_{ij}$ is:
$\hat{y}_{i}=\Theta\left(x_{ij}-c_{j}\right)$ (1)
$\hat{y}_{i}=1\rightarrow x_{ij}>c_{j}$ (2)
Here, $\Theta$ is the Heaviside step function, $c_{j}$ the cut’s boundary, and $\hat{y}_{i}$ the result. $\hat{y}_{i}=1$ means the event passes the cut $x_{ij}>c_{j}$ which is taken as a signal. We use the strictly greater than sign for clarity. This is a fundamental case where signal lies right of $c_{j}$ as a lower limit.
For the left case, with signal below the cut, the operation becomes:
$\hat{y}_{i}=\Theta\left(c_{j}-x_{ij}\right)$ (3)
$\hat{y}_{i}=1\rightarrow x_{ij}<c_{j}$ (4)
Here, $\hat{y}_{i}=1$ means the event passes the cut $x_{ij}<c_{j}$, taken as signal.
To make cuts learnable and auto-updated, we replace the non-differentiable $\Theta$ with the logistic function $\sigma_{L}$, a smooth approximation common neural networks:
$\sigma_{L}(z)=\frac{1}{1+e^{-z}}$ (5)
This enables us to define learnable cut operations for both cases:
$\hat{y}_{i}=\sigma_{L}\left(x_{ij}-b_{j}\right)$ (6)
$\hat{y}_{i}>t\rightarrow x_{ij}>\sigma_{L}^{-1}(t)+b_{j}$ (7)
$\hat{y}_{i}=\sigma_{L}\left(b_{j}-x_{ij}\right)$ (8)
$\hat{y}_{i}>t\rightarrow x_{ij}<b_{j}-\sigma_{L}^{-1}(t)$ (9)
Here, $b_{j}$ is a trainable bias, also the learnable cut value. $t$ is a probability threshold (often 0.5) to flag signal events.
To combine the left and right cases in a single operation, we introduce a trainable weight $w_{j}$:
$\hat{y}_{i}=\sigma_{L}\left(w_{j}x_{ij}-b_{j}\right)$ (10)
The cut’s behavior becomes:
$\hat{y}_{i}>t\rightarrow\begin{cases}x_{ij}>\text{boundary},&\text{if }\text{direction}>0\\ x_{ij}<\text{boundary},&\text{if }\text{direction}<0\end{cases}$ (11)
$\text{boundary}=\frac{1}{w_{j}}\left(\sigma_{L}^{-1}(t)+b_{j}\right)$ (12)
$\text{direction}=\text{sign}(w_{j})$ (13)
Here, the direction determines whether the signal lies above ($w_{j}>0$) or below ($w_{j}<0$) the boundary. This transforms a static cut to dynamic, learnable one, mimicking a neuron’s operation. Thus, training a neuron to tune $w_{j}$ and $b_{j}$ aligns with finding optimal boundary value—a clear connection from traditional cuts to neural networks.
Beyond single-sided cuts, double-sided cases—middle case (signal between two boundaries) and edge case (signal outside two boundaries)—are also prevalent. To manage these, we divide the observable’s distribution at $\text{center}_{j}$ (set manually from the distribution) into two ranges and then feed them into two learnable cuts:
$\hat{y}_{ij}^{\text{lower}}$ $=\sigma_{L}\left(w_{j}^{\text{lower}}x_{ij}-b_{j}^{\text{lower}}\right)$ (14)
$\hat{y}_{ij}^{\text{upper}}$ $=\sigma_{L}\left(w_{j}^{\text{upper}}x_{ij}-b_{j}^{\text{upper}}\right)$
Accordingly, the loss function shifts as follows. The loss $L_{ij}$ blends the masked average of both cut losses:
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}\operatorname{Mask}_{ij}^{\text{lower}}+L^{\text{upper}}_{ij}\operatorname{Mask}_{ij}^{\text{upper}}\right)$ (15)
$L_{ij}\left(y_{i},\hat{y}_{ij}\right)=-\left[y_{i}\log\left(\hat{y}_{ij}\right)+(1-y_{i})\log\left(1-\hat{y}_{ij}\right)\right]$ (16)
where $L$ is the binary cross-entropy loss between the true label $y_{i}$ and predicted output $\hat{y}_{ij}$. The masks are defined as:
$\text{Mask}^{\text{lower}}_{ij}=\begin{cases}1,&\text{if }x_{ij}<\text{center}_{j}\\ 0,&\text{otherwise}\end{cases}$ (17)
$\text{Mask}^{\text{upper}}_{ij}=\begin{cases}1,&\text{if }x_{ij}>\text{center}_{j}\\ 0,&\text{otherwise}\end{cases}$ (18)
To tie this together, we generalize all four cases (left, right, middle, and edge) using two sets of trainable parameters: $w_{j}^{\text{lower}}$, $b_{j}^{\text{lower}}$ for the lower range and $w_{j}^{\text{upper}}$, $b_{j}^{\text{upper}}$ for the upper range. These define a learnable cut for each observable, whether it’s a single-sided cut (left or right) or a double-sided one (middle or edge). For the simpler left and right cases, a single set of parameters can suffice, as shown in 11, but the two-set approach fully accommodates the more complex middle and edge scenarios. Based on later experimental results, we summarize the combination rules for the final learned cut:
  • The specified $\text{center}_{j}$ separates the lower cut and upper cut. A cut is invalid if its boundary crosses $\text{center}_{j}$; otherwise, it is valid.
  • If one cut is valid and the other invalid, the final cut follows the valid cut’s boundary and direction.
  • If both cuts are valid, we check their directions. If they align, the final cut is the intersection of both cuts’ ranges. If they differ, a positive lower ($w_{j}^{\text{lower }}>0$) and negative upper ($w_{j}^{\text{upper}}<0$) yield the middle case; otherwise, the edge case.
  • If both cuts are invalid, single-sided cuts (left or right) remain equivalent to their valid forms. Double-sided cuts are always taken as the edge case, forming the union of ranges to bypass observables with poor separation.
With this framework, we complete the transition from traditional cuts to fully learnable cuts.
2.2 Learnable importance
As the number of observables grows, determining their optimal combination becomes increasingly difficult. Unlike neural networks, which adapt to input features flexibly, traditional cut flows are sensitive to the order and selection of observables. Usually, highly-correlated observables are drop since they function similarly with each other. Those that have large overlap area between signal and background are also dropped since they don’t provide distinct characteristics between the two classes. We aim to let the LCF model to learn such an importance during the training process by providing a proper tool.
To achieve this, we introduce the Learnable Importance for each observable. For the $j$-th observable, we define a trainable parameter $s_{j}$, and the importance score $s^{\prime}_{j}$ is computed using the softmax function:
$s^{\prime}_{j}=\sigma_{S}(\mathbb{s})_{j}=\frac{e^{s_{j}}}{\sum_{k=1}^{F}e^{s_{k}}}$ (19)
Here $\mathbf{s}=\left[s_{1},s_{2},\ldots,s_{F}\right]$ is the vector of trainable parameters across all observables, and $\sigma_{S}$ denotes the softmax function. While softmax is not the only option—any differentiable function ensuring $\sum_{j=1}^{F}s_{j}^{\prime}=1$ could work—its prevalence in machine learning makes it a natural choice.
This importance score acts as a scaling factor for each observable’s contribution. For a data point $x_{ij}$, we define a scaled input:
$x_{ij}^{\prime}=s_{j}^{\prime}x_{ij}$ (20)
which feeds into the learnable cut operation:
$\hat{y}_{ij}=\sigma_{L}\left(w_{j}x_{ij}^{\prime}-b_{j}\right)$ (21)
Here, a higher $s_{j}^{\prime}$ retains a more complete range, emphasizing the correspondent observable’s role in signal-background separation, while a lower $s_{j}^{\prime}$ narrows it, reducing its influence when separation is poor due to highly-overlapped distributions.
To fully understand this mechanism, we analyze the gradient descent process, focusing on a single cut per observable for simplicity (noting that the two-cut setup from section 2.1 follows similarly). Using binary cross-entropy as the loss function, the total loss across $N$ events and $F$ observables is:
$L=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{F}L_{ij}$ (22)
where:
$L_{ij}=-[y_{i}\log(\hat{y}_{ij})+(1-y_{i})\log(1-\hat{y}_{ij})]$ (23)
$\hat{y}_{ij}=\sigma_{L}(z_{ij})$ (24)
$z_{ij}=w_{j}x^{\prime}_{ij}+b_{j}$ (25)
$x^{\prime}_{ij}=s^{\prime}_{j}x_{ij}$ (26)
$s^{\prime}_{j}=\sigma_{S}(\boldsymbol{s})_{j}$ (27)
The gradient of $L$ with respect to $s_{k}$ reveal how importance scores adjust during training. We compute:
$\frac{\partial L}{\partial s_{k}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{F}\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}\cdot\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}\cdot\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}}$ (28)
where:
$\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}=\frac{\hat{y}_{ij}-y_{i}}{\hat{y}_{ij}\left(1-\hat{y}_{ij}\right)}$ (29)
$\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}=\hat{y}_{ij}\left(1-\hat{y}_{ij}\right)$ (30)
$\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}=w_{j}x_{ij}$ (31)
Since $s^{\prime}_{j}$ depends on all $s_{k}$ via softmax, its gradient is a Jacobian matrix:
$\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=\begin{cases}s_{j}^{\prime}\left(1-s_{j}^{\prime}\right)&k=j\\ -s_{k}^{\prime}s_{j}^{\prime}&k\neq j\end{cases}$ (32)
Define:
$\delta_{j}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{ij}-y_{i}\right)w_{j}x_{ij}$ (33)
then:
$\frac{\partial L}{\partial s_{k}}=\sum_{j=1}^{F}\delta_{j}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=s_{k}^{\prime}\left(\delta_{k}-\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}\right)$ (34)
This gradient drives the learning dynamics. Here, $\delta_{k}$ reflects the $k$-th observable’s misalignment between predictions and labels, while $\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}$ is a weighted baseline across all observables. If $\delta_{k}>$ baseline, the gradient is positive, reducing $s_{k}$ and thus $s_{k}^{\prime}$. This shrink the value of the observable towards zero and the $s_{k}^{\prime}$ outside the parenthesis reduce its gradient making it rely more on other influential ones. If $\delta_{k}<$ baseline, the gradient is negative, increasing $s_{k}^{\prime}$, boosting a reliable observable to keep the full range of it. This process embeds importance directly into training. Such a learned importance offers insights into each observable’s role and plays as a potential threshold for dropping negligible ones.
2.3 Learnable cut flow
With the learnable cuts from section 2.1 and the learnable importance from section 2.2, we now build the LCF model on them, as shown in figure 1.
Figure 1: The structure of the learnable cut flow model.
The model takes the original data as inputs. Then data is normalized with the mean and standard deviation of the training data. Then data is split along the observable axis to feed into each learnable cut. The output of each learnable cut is collected at the end. When the LCF model performs inference on new data, a minimum importance ratio is applied, set to 0.05 by default. This means it automatically ignores cuts with observables below 5% of the average, where the average is defined as 1 divided by the number of features.
There are two strategies commonly adopted in the literature:
  • Parallel strategy: Each cut is optimized independently, focusing solely on the original data distribution to determine cut values, ignoring inter-observable correlations.
  • Sequential strategy: Cut values are optimized sequentially, with each cut adjusting based on the updated data distribution after applying the previous cut, fully considering correlations among different observables.
The parallel strategy is straightforward because it independently optimizes each cut, requiring all collected cut results to exceed a threshold at the end of the flow for an event to pass all cuts during optimization.
In the sequential strategy, we cannot change the input data shape during forward propagation, as neurons must preserve the original structure for stacking. Instead, we mask events by setting their contribution to the loss function to zero, effectively excluding them without altering the data shape. We modify the loss per event $i$ and observable $j$ for the sequential strategy as:
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}+L^{\text{upper}}_{ij}\right)\cdot\text{Mask}_{i,j-1}$ (35)
where $L_{ij}^{\text{lower}}$ and $L_{ij}^{\text{upper}}$ are the binary cross-entropy losses for the two ranges (as in section 2.1), and $\text{Mask}_{i,j-1}$ is the mask from the previous cut, defined as:
$\text{Mask}_{i,j}=\begin{cases}1&\text{if }j=0\\ \Theta(\hat{y}^{\text{lower}}_{i,j-1}-t)\cdot\Theta(\hat{y}^{\text{upper}}_{i,j-1}-t)&\text{if }j>0\end{cases}$ (36)
The mask takes the intersection of the outputs from both ranges, ensuring the loss accounts for the effect of the previous cut’s decisions.
With these parallel and sequential strategies, we’ve constructed a learnable cut flow model that fully mimics the traditional approach, enhancing it with the adaptability and interpretability of neural networks.
3 Experiment setup
3.1 Datasets
This study employs two types of datasets: a series of synthetic mock datasets to demonstrate the effectiveness of the proposed neural network under controlled and simplified conditions, and a real dataset from 1603.09349 to benchmark its performance against baseline models.
The synthetic datasets are constructed from different combinations of ten features sampled from distinct normal distributions ($N\left(\mu,\sigma^{2}\right)$ where $\mu$ is mean and $\sigma$ is the standard deviation). The parameters are specified in table 1 and distributions are presented in figure 2.
Feature Signal Background Comment
$x_{1}$ $N(-2,4)$ $N(2,4)$ Left case
$x_{2}$ $N(2,4)$ $N(-2,4)$ Right case
$x_{3}$ $N(0,4)$ $N(5,4)+N(-5,4)$ Middle case
$x_{4}$ $N(5,4)+N(-5,4)$ $N(0,4)$ Edge case
$x_{5}$ $N(-1,4)$ $N(1,4)$ Weakly separated
$x_{6}$ $N(-3,4)$ $N(3,4)$ Strongly separated
$x_{7}$ $N(-5,4)$ $N(-5,4)$ Redundant
$x_{8}$ $N(5,4)$ $N(5,4)$ Redundant
$x_{9}$ $0.9\times x_{1}^{\text{signal }}+N(0,1)$ $0.9\times x_{1}^{\text{background }}+N(0,1)$ Highly correlated
$x_{10}$ $0.7\times x_{1}^{\text{signal }}+N(0,1)$ $0.7\times x_{1}^{\text{background }}+N(0,1)$ Highly correlated
Table 1: Parameters of signal and background features for synthetic mock datasets.
Figure 2: Mock feature distributions.
Each dataset is designed for a specific purpose and with correlations shown in figure 3:
  • Mock1 consists of four features ($x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$) to validate whether the learned decision boundaries are reasonable and intuitive across four fundamental cases: signals located at the left, right, middle, and edge.
  • Mock2 comprises three features ($x_{1}$, $x_{5}$, $x_{6}$) to assess the interpretability of learned feature importance by adjusting the signal-background overlap in $x_{1}$, with $x_{5}$ reducing the overlap and $x_{6}$ increasing it.
  • Mock3 includes two redundant features ($x_{7}$, $x_{8}$) alongside $x_{1}$ to assess the model’s robustness to irrelevant inputs.
  • Mock4 combines $x_{1}$ with two highly correlated features ($x_{9}$, $x_{10}$) to simulate a common challenge in phenomenology data analysis.
  • Mock5 incorporates a typical set of features ($x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$, $x_{5}$, $x_{7}$, $x_{9}$) to represent a realistic dataset.
  • Mock6 randomly reorders the features of Mock5 to evaluate the model’s sensitivity to feature order.
Figure 3: Feature correlations of mock datasets.
The real dataset was created using standard simulation tools in high energy physics. The signal, diboson production $pp\rightarrow W^{+}W^{-}\rightarrow qqqq$, produces two fat jets exhibiting a 2-prong substructure. In contrast, the background, QCD dijet production $pp\rightarrow qq,qg,gg$ predominantly yields 1-prong jets.
Events were generated using MadGraph 5 v2.2.3 for hard scattering, Pythia v6.426 for showering and hadronization, and Delphes v3.2.0 for detector response. Jets were reconstructed with the anti-$k_{t}$ algorithm and parameter $R=1.2$ using FastJet v3.1.2. Their constituents were then reclustered with the $k_{t}$ algorithm and parameter $R=0.2$ to form subjets. Subjets with $p_{T}^{\text{sub}}<p_{T}^{\text{jet}}\times 3\%$ were discarded and the remaining subjets were used to construct the trimmed jets. To ensure the collinear $W$ bosons decays, only jets with $p_{T}^{\text{trim}}\in[300,400]$ GeV were retained.
Six high-level observables were calculated: the invariant mass of the trimmed jet ($M_{\text{jet}}$), the N-subjettiness ratio ($\tau_{21}^{\beta=1}$), and four ratios of energy correlation functions ($C_{2}^{\beta=1},C_{2}^{\beta=2},D_{2}^{\beta=1},D_{2}^{\beta=2}$). Their distributions are shown in figure 4 and their correlations in figure 5.
Figure 4: Distributions of jet substructure variables. Top row: jet mass (left) and $\tau_{21}$ with $\beta=1$ (right). Middle row: $C_{2}$ with $\beta=1$ (left) and $\beta=2$ (right). Bottom row: $D_{2}$ with $\beta=1$ (left) and $\beta=2$ (right).
Figure 5: Observable correlations of the diboson dataset.
Both datasets contains 200,000 samples evenly divided between signal and background. The data are split into a training set (50%) and a test set (50%). To maintain computational efficiency, samples of the real dataset are randomly drawn from the original non-pile up test set. Furthermore, the original observable distributions show a significant presence of low values for all observables except $M_{\text{jet}}$. To mitigate outliers and ensure stable training, only samples within the 5th to 95th percentile range of all observables are retained. Both dataset types define a binary classification task, allowing the learnable cut flow models to be well trained and provide insights into the neural network’s decision-making process.
3.2 Baseline models
Two baseline models are implemented: a boosted decision tree (BDT) and a multilayer perceptron (MLP):
  • The BDT, built with Scikit-Learn, consists of 100 estimators with a maximum depth of 3. It is trained directly on raw input data without preprocessing as it is insensitive to feature scaling. The output represents the probability of the signal event.
  • The MLP, built with Keras, includes a standard normalization layer for input preprocessing and five hidden layers with 16, 32, 64, 32, and 16 units respectively. All hidden layers use the rectified linear unit (ReLU) activation function. The output layer applies a sigmoid function to produce the signal probability.
All models involved (BDT, MLP, LCF in two strategies) are trained using the same configuration: binary cross entropy as the loss function, Adam as the optimizer with a learning rate 0.001, a batch size of 512, and 200 training epochs. Early stopping is not applied as training of LCF exhibits plateau phases in loss. The centers for LCF model are set as -2, 2, 0, 0, -1, -3, -5, 5, -1.8, -1.4 for each mock feature, and 80, 0.15, 0.025, 2, 2, 0.3 for each observable in the diboson dataset.
4 Results
4.1 Learned cuts
To validate the ability of LCF models to learn optimal cut positions, we evaluate their performance on the Mock1 dataset, which includes four features illustrating fundamental cut scenarios: left, right, middle, and edge cases. The learned cuts are shown alongside signal and background distributions in figure 6 (parallel LCF) and figure 7 (sequential LCF).
In the parallel strategy, the model analyzes each feature’s distribution independently to optimize its cut positions. As shown in figure 6, the resulting cuts effectively isolate signal-rich regions: $x_{1}$ captures the left tail of the signal, $x_{2}$ the right tail, $x_{3}$ the central peak, and $x_{4}$ the outer tails. These positions align well with the boundaries designed for Mock1. By contrast, the sequential strategy applies cuts iteratively, modifying the distribution after each cut to account for inter-feature correlations. Figure 7 illustrates how the cut positions adapt to these updates, with subsequent cuts refining the remaining signal region to improve signal-background separation as the feature space is progressively filtered by earlier cuts.
The importance scores shown in figure 9 and figure 9 reflect how the LCF prioritizes each feature. In the parallel strategy, scores are relatively balanced: features with single-sided cuts ($x_{1}$ and $x_{2}$) have scores 0.179 and 0.185, respectively, compared to features with double-sided cuts ($x_{3}$ and $x_{4}$) at 0.135 and 0.321. By contrast, the importance shifts in the sequential strategy as prior cuts filter out a fraction of background events: $x_{1}$ increases to 0.224 while $x_{2}$ decreases to 0.164; similarly, $x_{3}$ increases to 0.351 and $x_{4}$ decreases to 0.260.
The learned cuts and importance scores demonstrate that LCF models can replicate the traditional cut-based methods while also dynamically adapting to the evolving feature space during training.
Figure 6: Learned cuts from the parallel LCF on the Mock1 dataset.
Figure 7: Learned cuts from the sequential LCF on the Mock1 dataset.
Figure 8: Learned importance from the parallel LCF on the Mock1 dataset.
Figure 9: Learned importance from the sequential LCF on the Mock1 dataset.
4.2 Learned importance
To further evaluate the LCF model’s ability to prioritize features, we analyze the importance scores learned on the Mock2 dataset, which includes three features ($x_{1}$, $x_{5}$, $x_{6}$) with regular, weak, and strong signal-background separation. The signal and background distributions, along with the learned cuts, are shown in figure 10 (parallel LCF) and figure 11 (sequential LCF). The corresponding importance scores are shown in figure 13 and figure 13.
In the parallel strategy, the LCF model assigns importance scores that closely match each feature’s intended role: compared to $x_{1}$’s regular signal-background separation, $x_{5}$ decrease it while $x_{6}$ increase it. In the sequential strategy, however, cuts modify the features space step by step, which reduces $x_{5}$’s effective separation and thus lowers its importance. Meanwhile, $x_{6}$ maintains a clear boundary between signal and background, resulting in a score similar to $x_{1}$.
These results demonstrate that the learnable importance mechanism captures each features’ discriminative power and adapts dynamically under both the parallel and sequential strategies.
Figure 10: Learned cuts from the parallel LCF on the Mock2 dataset.
Figure 11: Learned cuts from the sequential LCF on the Mock2 dataset.
Figure 12: Learned importance from the parallel LCF on the Mock2 dataset.
Figure 13: Learned importance from the sequential LCF on the Mock2 dataset.
4.3 Robustness to redundancy
To assess how robust the LCF models are when presented with redundant features, we train them on the Mock3 dataset. In this setup, $x_{1}$ provides a clear signal-background separation, whereas $x_{7}$ and $x_{8}$ offer little to no additional discriminative power. As shown in figure 14, the learned cuts for $x_{7}$ and $x_{8}$ appear almost randomly placed, reflecting their limited ability to distinguish signal and background. The corresponding importance scores in figure 17 confirm that the model recognizes these two features contribute minimally to separation.
In the sequential strategy, the first cut leads to a nearly complete overlap of signal and background for $x_{7}$ and $x_{8}$ as shown in figure 15. This, in turn, results in almost zero importance for these features (see figure 17).
These findings highlight the LCF models’ robustness in detecting and effectively ignoring features with negligible separation power.
Figure 14: Learned cuts from the parallel LCF on the Mock3 dataset.
Figure 15: Learned cuts from the sequential LCF on the Mock3 dataset.
Figure 16: Learned importance from the parallel LCF on the Mock3 dataset.
Figure 17: Learned importance from the sequential LCF on the Mock3 dataset.
4.4 Robustness to high correlation
High correlation between observables implies that a cut on one feature effectively constrains the others, posing a challenge for cut-based methods. To evaluate the LCF models’ robustness in this scenario, we train them on the Mock4 dataset which includes $x_{1}$ alongside highly correlated features $x_{9}$ and $x_{10}$. The resulting learned cuts are shown in figure 18 (parallel LCF) and figure 19 (sequential LCF), with corresponding importance scores shown in figure 21 and figure 21.
Figure 18: Learned cuts from the parallel LCF on the Mock4 dataset.
Figure 19: Learned cuts from the sequential LCF on the Mock4 dataset.
Figure 20: Learned importance from the parallel LCF on the Mock4 dataset.
Figure 21: Learned importance from the sequential LCF on the Mock4 dataset.
In the parallel strategy, the LCF model treats each feature independently, so the high correlation doesn’t significantly affect the learned cuts or importance scores. These scores remain consistent with the dataset’s design. By contrast, since the first cut in the sequential strategy reshapes the subsequent feature distributions, $x_{9}$ and $x_{10}$ end up with lower importance.
Both strategies successfully identify the proper cut boundaries under high correlation, demonstrating that the LCF models remain robust in this setting.
4.5 Generalization across feature sets
To evaluate the generalization ability, we train the LCF models on the Mock5 and Mock6 datasets. Mock5 consists of a typical combination of features, and Mock6 introduces random reordering of features to test whether the model remains robust to different feature arrangements. The learned cuts for Mock5 are shown in figure 22 (parallel LCF) and figure 23 (sequential LCF) with corresponding learned importance scores shown in figure 25 and figure 25. The figures related to Mock6 are 22, 23, 25, and 25..
Figure 22: Learned cuts from the parallel LCF on the Mock5 dataset.
Figure 23: Learned cuts from the sequential LCF on the Mock5 dataset.
Figure 24: Learned importance from the parallel LCF on the Mock5 dataset.
Figure 25: Learned importance from the sequential LCF on the Mock5 dataset.
Figure 26: Learned cuts from the parallel LCF on the Mock6 dataset.
Figure 27: Learned cuts from the sequential LCF on the Mock6 dataset.
Figure 28: Learned importance from the parallel LCF on the Mock6 dataset.
Figure 29: Learned importance from the sequential LCF on the Mock6 dataset.
In the parallel strategy for Mock5, the learned cuts are consistent with those observed in previous Mock datasets except for feature $x_{7}$ whose cut covers the full range due to its lack of signal-background separation as shown in figure 22. Unlike the seemingly random cuts for $x_{7}$ in section 4.3, this result more clearly demonstrates that the model directly ignores such a feature. This matches its importance score which is zero in figure 25. The relative importance among $x_{1}$ to $x_{4}$ remains similar to Mock1. Feature $x_{9}$ is more important than $x_{5}$, which aligns with their signal-background separation. In the sequential strategy, it clearly shows that the cuts for $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ effectively separate signal from background. By contrary, $x_{5}$’s cut has limited separation and $x_{7}$ and $x_{9}$ cover the full range, which matches their zero importance scores in figure 25.
The results from the reordered dataset Mock6 are interesting. In the parallel strategy, since every feature is treated separately, the importance scores are nearly identical to those in Mock5 (figure 29). By contrast, in the sequential strategy, the presence of a less informative feature like $x_{7}$ early in the sequence doesn’t prevent the model from converging correctly. The cut boundaries for $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ remain clear and effective while $x_{7}$ and $x_{9}$ cover the full range as in Mock5. The cut for $x_{5}$ appears less constrained than in Mock5, leading to a small non-zero importance in figure 29, though its separation remains limited.
These findings show that our models are robust when encountering all typical features: informative, redundant, highly correlated, and robust to feature ordering.
4.6 Benchmark on diboson classification
Finally, we evaluate the LCF model on the real diboson dataset in both strategies and also compare its performance with other baseline models. The learned cuts are shown in figure 30 (parallel LCF) and figure 31 (sequential LCF). The corresponding learned importance scores are shown in figure 33 and figure 33. The performance is summarized in table 2. The table presents true positive (TP, the number of retained signal events), false positive (FP, the number of inevitable background events), accuracy (the number of correctly tagged events), precision (the ratio of correctly tagged signal), and significance which is calculated by $\text{TP}/\sqrt{\text{FP}}$.
Figure 30: Learned cuts from the parallel LCF on the diboson dataset.
Figure 31: Learned cuts from the sequential LCF on the diboson dataset.
Figure 32: Learned importance from the parallel LCF on the diboson dataset.
Figure 33: Learned importance from the sequential LCF on the diboson dataset.
Model TP FP Accuracy Precision Significance
BDT 41213 6245 0.8754 0.8684 521.5164
MLP 41135 5760 0.8808 0.8772 542.0012
LCF (Parallel) 24230 2849 0.6953 0.8948 453.9492
LCF (Sequential) 40520 6710 0.8601 0.8579 494.6616
Table 2: Performance metrics of different models on the diboson dataset.
In the parallel strategy, the learned cuts are intuitively placed at the boundaries where signal and background distributions diverge, effectively separating the two classes across all observables. In contrast, in the sequential strategy, the first two cuts on $M_{\mathrm{jet}}$ and $C_{2}^{\beta=1}$ filter out most background events. This causes the signal and background to almost completely overlap in the distributions of the later observables ($\tau_{21}^{\beta=1}$, $C_{2}^{\beta=2}$, $D_{2}^{\beta=1}$, $D_{2}^{\beta=2}$), making their cuts less effective. To confirm whether these overlapped observables are truly negligible, we examine the learned importance scores. In both strategies, all observables are assigned fair importance scores, with none close to zero, indicating that there are still signal samples in the long-tail regions of these distributions. Comparing the scores, $D_{2}^{\beta=2}$ ranks second in the parallel strategy and is more important than $D_{2}^{\beta=1}$, which aligns with the smaller overlap area in its distribution. In the sequential strategy, this trend persists, though the importance of later observables is generally reduced. Notably, $\tau_{21}^{\beta=1}$ is less important than both ${C_{2}^{\beta=1}}^{\beta=1}$ and $C_{2}^{\beta=2}$ in both strategies. This is likely because importance is more related to the absolute area of separation as observed in the figures.
Among all the tested models, the LCF model, as a cut-based method, exhibits the highest precision despite a lower accuracy. The parallel strategy effectively minimizes background events to a small fraction, albeit at the expense of a greater signal loss compared to other models. In contrast, the sequential strategy behave more like a classifier, enhancing accuracy by systematically checking each distribution while considering the effects of prior cuts.
We further investigate performance differences through two additional experiments. The first experiment only keeps the observables with the initial importance scores above the average (1/7) namely $M_{\text{jet}}$, $D_{2}^{\beta=1}$, and $D_{2}^{\beta=2}$. The second experiment excludes highly-correlated observables, retaining $M_{\text{jet}}$, $C_{2}^{\beta=2}$, $D_{2}^{\beta=2}$, and $\tau_{21}^{\beta=1}$, as informed by the correlation figure. The results of these experiments are shown in table 3 and table 4.
Model TP FP Accuracy Precision Significance
BDT 42046 7422 0.8567 0.8500 488.0499
MLP 42408 7532 0.8597 0.8492 488.6441
LCF (Parallel) 34266 4125 0.8029 0.8926 533.5209
LCF (Sequential) 40027 7078 0.8366 0.8497 475.7708
Table 3: Performance metrics of different models on the first dataset.
Model TP FP Accuracy Precision Significance
BDT 41066 6874 0.8655 0.8566 495.3106
MLP 40915 6568 0.8675 0.8617 504.8542
LCF (Parallel) 28111 3751 0.7376 0.8823 458.9895
LCF (Sequential) 40611 7304 0.8540 0.8476 475.1856
Table 4: Performance metrics of different models on the second dataset.
Retaining observables with higher-than-average importance leads to clear improvements: the parallel LCF model achieves a significance of 533 in table 3 (up from 453.95 in table 2), surpassing the 488 of both the BDT and MLP. In contrast, removing highly-correlated observables has little effect. This prove once again that the learnable importance is not only a metrics but also a practical standard to select observables.
5 Conclusion
In this work, we introduce the Learnable Cut Flow (LCF), a novel neural network framework that bridge the interpretability of traditional cut-based methods with the adaptability of modern machine learning techniques in high-energy physics. By transforming the manual process of cut selection into a fully differentiable, data-driven operation, LCF offers a transparent and actionable alternative to opaque black-box models. Two widely-used strategies are also implemented via mask operations making LCF a fully interpretable model. We also propose the Learnable Importance to dynamically adjust each observable’s contribution to the final classification results.
Testing on six synthetic mock datasets and a realistic diboson vs. QCD dataset demonstrated LCF’s versatility and effectiveness. The model accurately identifies optimal cut boundaries across a range of feature distributions, including left, right, middle, and edge cases, while dynamically adapting to varying degrees of signal-background separation. The Learnable Importance mechanism successfully assigns higher weights to features with stronger discriminative power, effectively suppresses redundant or highly correlated observables, and remains robust to feature ordering. On the diboson dataset, LCF initially lagged behind boosted decision trees (BDT) and multilayer perceptrons (MLP) in raw performance. However, by pruning less critical observables guided by learned importance, the parallel LCF achieved a significance of 533.52, surpassing the baselines (488.05 for BDT and 488.64 for MLP), underscoring the practical utility of this metric.
LCF’s ability to emulate human cut-searching strategies while embedding them within a trainable framework marks a significant step forward. It retains the simplicity and physicist-friendly insights of traditional cut flows, yet harnesses the power of neural networks to handle complex, high-dimensional data. This hybrid approach not only enhances classification performance but also provides interpretable outputs—cut boundaries and importance scores—that can guide observable selection in real-world analyses.
Acknowledgements.
Hao Sun is supported by the National Natural Science Foundation of China (Grant No. 12075043, No. 12147205).